# On the distance problem over finite p-adic rings

### v2

@@ 4.1 Extension theorems
Let $r\geq 1$ be an integer. Let $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}$. Then
@@
Let $p\equiv 1\,(\textsf{mod }4)$ and $r\geq 1$ be an integer. Let $\mathbf{m}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ be such that $\mathbf{m}\neq\mathbf{0}$. Denote by $d\sigma_{r}$ the normalized surface measure on $\textsf{orb}_{r}(\mathbf{m})$. Write $\mathbf{m}=p^{v_{m}}\tilde{\mathbf{m}}$ with $\tilde{\mathbf{m}}\in(\mathbb{Z}/p^{r-v_{m}}\mathbb{Z})^{2}$ and $\|\tilde{\mathbf{m}}\|\not\equiv 0\,(\textsf{mod }p)$. Then
@@ A $F$-sphere centered at $x$ of radius $j$ is the set
With Theorem 6.3 in hand, Theorem 1.14 now follows from a combinatorial argument which is identical to [22, Proof of Theorem 3.1] with Theorem 6.3 in place of [22, Lemma 2.2]. So we omit the details.

### v3

@@ 4.1 Extension theorems
Let $p$ be an odd prime, and $r\geq 1$ be an integer. Let $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}$. Then
@@
Let $p\equiv 1\,(\textsf{mod }4)$ and $r\geq 1$ be an integer. Let $\mathbf{m}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ be such that $\mathbf{m}\neq\mathbf{0}$. Denote by $d\sigma_{r}$ the normalized surface measure on $\textsf{orb}_{r}(\mathbf{m})$. Suppose that $\mathbf{m}=p^{v_{m}}\tilde{\mathbf{m}}$ with $\tilde{\mathbf{m}}\in(\mathbb{Z}/p^{r-v_{m}}\mathbb{Z})^{2}$ and $\|\tilde{\mathbf{m}}\|\not\equiv 0\,(\textsf{mod }p)$. Then
@@ A $F$-sphere centered at $x$ of radius $j$ is the set
With Theorem 6.3 in hand, Theorem 1.14 now follows from a combinatorial argument which is identical to [23, Proof of Theorem 3.1] with Theorem 6.3 in place of [23, Lemma 2.2]. So we omit the details.